# Two prover perfect zero knowledge for MIP*

### v1

@@ 8. Perfect zero knowledge
Given a BCS $B=(X,\{(V_{i},C_{i})\}^{m}_{i=1})$ and $n\geq 1$, let $Y=X\times[n]$, and $W_{i}=V_{i}\times[n]$ for any $1\leq i\leq m$. To make the elements of $Y$ look more like variables, we denote $(x,i)$ by $x(i)$. Let $D_{i}\subseteq\mathbb{Z}_{2}^{W_{i}}$ be the set of assignments $\phi$ to $W_{i}$ such that the assignment $\psi$ to $V_{i}$ defined by $\psi(x)=\psi(x(1))\cdots\psi(x(n))$ is in $C_{i}$. The obliviation of $B$ of degree $n$ is the constraint system $\operatorname{Obl}_{n}(B)=(Y,\{W_{i},D_{i}\}_{i=1}^{m})$.
@@
Suppose $B=(X,\{(V_{i},C_{i})\}^{m}_{i=1})$ is a BCS, and let $B^{\prime}=\operatorname{Obl}_{n}(B)$ for some $n\geq 1$. Then there is a classical homomorphism $\alpha:\mathcal{A}(B)\to\mathcal{A}(B^{\prime})$ such that $\alpha(\sigma_{i}(x))=\sigma_{i}(x(1)\cdots x(n))$ for all $i\in[m]$ and $x\in V_{i}$, where $\sigma_{i}$ is in the inclusion of the $i$th factor for $\mathcal{A}(B)$ and $\mathcal{A}(B^{\prime})$.
Furthermore, if $\pi$ is a probability distribution on $[m]\times[m]$, and $\tau$ is a tracial state on $\mathcal{A}(B)$, then there is a tracial state $\widetilde{\tau}$ on $\mathcal{A}(B^{\prime})$ such that $\tau=\widetilde{\tau}\circ\alpha$, $\operatorname{def}(\widetilde{\tau};\mu_{\pi})=\operatorname{def}(\tau;\mu_{\pi})$, and $\widetilde{\tau}(\prod_{(i,x)\in S}\sigma_{i}(x))=0$ for any ordered set $S$ of pairs $(i,x)$ with $x\in W_{i}$ and $11\leq|S|\leq n-$ 1.
@@
Define $f_{i}:\mathbb{Z}_{2}^{W_{i}}\to\mathbb{Z}_{2}^{V_{i}}$ for each $i\in[m]$ by $f_{i}(\phi)(x)=\phi(x(1))\cdots\phi(x(n))$ for $\phi\in\mathbb{Z}_{2}^{W_{i}}$ and $x\in V_{i}$. By definition, $\phi\in D_{i}$ if and only if $f_{i}(\phi)\in C_{i}$, so $f_{i}(D_{i})=C_{i}$. If $f_{i}(\phi)(x)\neq f_{j}(\psi)(x)$ for some $\phi\in\mathbb{Z}_{2}^{W_{i}}$, $\psi\in\mathbb{Z}_{2}^{W_{j}}$, and $x\in V_{i}\cap V_{j}$, then we must have $\phi(x(i))\neq\psi(x(i))$ for some $i$. Since
$\sigma_{i}(x(1)\cdots x(n))=\sum_{\phi\in\mathbb{Z}_{2}^{W_{i}}}f_{i}(\phi)(x)\Phi_{W_{i},\phi}$
for all $x\in V_{i}$, $i\in[m]$, the functions $f_{i}$ correspond to a classical homomorphism $\alpha:\mathcal{A}(B)\to\mathcal{A}(B^{\prime})$ with $\alpha(\sigma_{i}(x))=\sigma_{i}(x(1)\cdots x(n))$ for all $i\in[m]$ and $x\in V_{i}$.
Conversely, given $y\in\mathbb{Z}_{2}^{X\times[n-1]}$ and $\phi\in\mathbb{Z}_{2}^{V_{i}}$, define $\phi_{y}\in\mathbb{Z}_{2}^{W_{i}}$ by $\phi_{y}(x(1))=\phi(x)y(x,1)$, $\phi_{y}(x(j))=y(x,j-1)y(x,j)$ for $12\leq j\leq n-$ 1, and $\phi_{y}(x(n))=y(x,n-1)$. Since $f_{i}(\phi_{y})=\phi$, the function $\phi\mapsto\phi_{y}$ sends $C_{i}$ to $D_{i}$. Also if $\phi\in\mathbb{Z}_{2}^{V_{i}}$ and $\psi\in\mathbb{Z}_{2}^{V_{j}}$, then $\phi_{y}|_{W_{i}\cap W_{j}}\neq\psi_{y}|_{W_{i}\cap W_{j}}$ if and only if $\phi|_{V_{i}\cap V_{j}}\neq\psi|_{V_{i}\cap V_{j}}$, so the functions $\phi\mapsto\phi_{y}$ determine a local homomorphism $\beta_{y}:\mathcal{A}(B^{\prime})\to\mathcal{A}(B)$ with $\beta_{y}(\sigma_{i}(x(1)))=\sigma_{i}(x)y(x,1)$, $\beta_{y}(\sigma_{i}(x(j)))=y(x,j-1)y(x,j)$ for $12\leq j\leq n-$ 1, and $\beta_{y}(\sigma_{i}(x(n)))=y(x,n-1)$ for all $i\in[m]$ and $x\in V_{i}$.
Define a tracial state $\widetilde{\tau}$ on $\mathcal{A}(B^{\prime})$ by $\widetilde{\tau}=2^{-|X|(n-1)}\sum_{y}\tau\circ\beta_{y}$, where the sum is over all $y\in\mathbb{Z}_{2}^{X\times[n-1]}$. Since $\beta_{y}\circ\alpha$ is the identity on $\mathcal{A}(B)$, $\widetilde{\tau}\circ\alpha=\tau$. Since $\beta_{y}$ and $\alpha$ are $1$-homomorphisms,
@@
Finally, if $S$ is an ordered set of pairs $(i,x)$ with $x\in W_{i}$, then there is an element $a\in\mathcal{A}(B)$ and set $S^{\prime}\subseteq X\times[n-1]$ such that
$\beta_{y}(\prod_{(i,x)\in S}\sigma_{i}(x))=m_{y}\tau(a)$
for all $y\in\mathbb{Z}_{2}^{X\times[n-1]}$, where $m_{y}:=\prod_{(x,j)\in S^{\prime}}y(x,j)$. If $|S|\leq n-1$, then $S^{\prime}$ is non-empty. Hence
$\widetilde{\tau}(\prod_{(i,x)\in S}\sigma_{i}(x))=2^{-|X|(n-1)}\sum_{y}m_{y}\tau(a),$
and $\sum_{y}m_{y}=0$ if $S^{\prime}$ is non-empty. ∎
@@
$Y=X\sqcup\{T_{i}(p,q),r_{i}(j,k):(i,p,q,k,j)\in{[m]\times[3]\times[d_{i}]\times[2]\times[d_{i}-1]}\}$
@@
As mentioned above, the product in the constraints on the permutation-valued variables in parts (1)-(4) of the definition is the group product in $S_{5}$. The constraints in part (1) involve both original variables $x_{p}$ and permutation-valued variables $T_{i}(1,p)$, and say that the value of $T_{i}(1,q)$ is either $\pi^{(iq)}_{1}$ or $\pi^{(iq)}_{-1}$ depending on the value of $x_{q}$. In part (4), $x$ and $y$ can be either an original or a permutation-valued variable. If one of them is a permutation-valued variable, then all the corresponding boolean variables encoding the permutation-valued variable are included in the constraint (so the constraint on $x$ and $y$ may involve up to $14$ boolean variables). Since the constraints in part (4) are trivial, they do not contribute to $D_{i}$, but they are included in the list of clauses $(W_{ij},D_{ij})$ of the subdivided tableau. The point of the constraints in part (4) is that, with them, $\operatorname{Tab}_{sub}(B)$ is a subdivision of $\operatorname{Tab}(B)$. Finally, observe that the constraints $D_{i}$ encode the constraints $C_{i}$ as follows:
@@
Suppose $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, and let $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$. If $\psi\in D_{i}$, then $\psi|_{V_{i}}\in C_{i}$. Conversely, if $r\in\S_{5}^{R_{i}}$, where $R_{i}=\{r_{i}(j,k):(j,k)\in[2]\times[d_{i}-1]\}$ is the set of randomizers in $W_{i}$, and $\phi\in C_{i}$, then there is a unique element $\phi_{r}\in D_{i}$ such that $\phi_{r}|_{V_{i}}=\phi$ and $\phi_{r}|_{R_{i}}=r$.
In this lemma, the statement that $\phi_{r}|_{R_{i}}=r$ means that for every randomizer $r_{i}(j,k)\in R_{i}$, the restriction of $\phi$ to the boolean variables corresponding to $r_{i}(j,k)$ is the encoding of the permutation $r(r_{i}(j,k))$. Although the permutation-valued variables in $\operatorname{Tab}(B)$ are shorthand for boolean variables, it is helpful to be able to work with the permutation-valued variables directly in $\mathcal{A}(\operatorname{Tab}(B))$. Suppose for a moment that $x_{1},\ldots,x_{7}$ are variables in a set $V$, and $C$ is a constraint on $V$ which includes the requirement that $x_{1},\ldots,x_{7}$ encode a permutation-valued variable $x$. Let $S=\{x_{1},\ldots,x_{7}\}$. If $\phi\in\mathbb{Z}_{2}^{S}$, then $\Phi_{S,\phi}=0$ in $\mathcal{A}(V,C)$ unless $\phi$ is the binary representation of an index $1200\leq j<$ 120, in which case we also write $\Phi_{S,\phi}$ as $\Phi_{S,j}$. Hence the subalgebra of $\mathcal{A}(V,C)$ is generated by the single unitary $\sum_{j=0}^{119}e^{2\pi ij/120}\Phi_{S,j}$, which we denote by the same symbol as the permutation-valued variable $x$. In particular, if $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ and $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$ as in Definition 8.6, then we can refer to $T_{i}(p,q)$ and $r_{i}(j,k)$ as unitary elements of $\mathcal{A}(W_{i},D_{i})$ of order $120$, and they generate the same subalgebra as the boolean variables encoding them. Since these variables do not occur in any other context $W_{j}$ for $j\neq i$, we also use $T_{i}(p,q)$ and $r_{i}(j,k)$ to refer to $\sigma_{i}(T_{i}(p,q))$ and $\sigma_{i}(r_{i}(j,k))$ in $\mathcal{A}(\operatorname{Tab}(B))$. We use the same convention for $\mathcal{A}(\operatorname{Tab}_{sub}(B))$.
The algebra $\mathcal{A}(\operatorname{Tab}(B))$ is generated by the original variables and the randomizers.
@@
Suppose $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, and let $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$. Let $R_{i}=\{r_{i}(j,k):(j,k)\in[2]\times[d_{i}-1]\}$ be the set of randomizers in $W_{i}$, and let $R=\bigcup_{i}R_{i}$. Then $\mathcal{A}(W_{i},D_{i})$ is generated as an algebra by $V_{i}\cup R_{i}$, and $\mathcal{A}(\operatorname{Tab}(B))$ is generated by $\bigcup_{i}\{\sigma_{i}(x):x\in V_{i}\}\cup R$.
This means that a homomorphism $\mathcal{A}(\operatorname{Tab}(B))\to\mathcal{A}(B)$ is completely described by its action on $V_{i}\cup R$. The following lemma extends Lemma 8.7 to weighted BCS algebras.

### v2

@@ 9. Perfect zero knowledge
Given a BCS $B=(X,\{(V_{i},C_{i})\}^{m}_{i=1})$ and $n\geq 1$, let $Z=X\times[n]$, and $U_{i}=V_{i}\times[n]$ for any $1\leq i\leq m$. To make the elements of $Z$ look more like variables, we denote $(x,i)$ by $x(i)$. Let $E_{i}\subseteq\mathbb{Z}_{2}^{U_{i}}$ be the set of assignments $\phi$ to $U_{i}$ such that the assignment $\psi$ to $V_{i}$ defined by $\psi(x)=\psi(x(1))\cdots\psi(x(n))$ is in $C_{i}$. The obliviation of $B$ of degree $n$ is the constraint system $\operatorname{Obl}_{n}(B)=(Z,\{(U_{i},E_{i})\}_{i=1}^{m})$.
@@
Suppose $B=(X,\{(V_{i},C_{i})\}^{m}_{i=1})$ is a BCS, and let $B^{\prime}=\operatorname{Obl}_{n}(B)$ for some $n\geq 1$, using the notation from Definition 9.3. Then:
  1. (a)
    There is a classical homomorphism $\alpha:\mathcal{A}(B)\to\mathcal{A}(B^{\prime})$ such that $\alpha(\sigma_{i}(x))=\sigma_{i}(x(1)\cdots x(n))$ for all $i\in[m]$ and $x\in V_{i}$, where $\sigma_{i}$ is the inclusion of the $i$th factor for $\mathcal{A}(B)$ and $\mathcal{A}(B^{\prime})$.
  2. (b)
    Let $\Gamma$ be the set of sequences $x_{1},\ldots,x_{k}$ in $Z$ of length $11\leq k\leq n-$ 1, such that there is some $i\in[k]$ with $x_{i}\neq x_{j}$ for all $j\in[k]\setminus\{i\}$. If $\pi$ is a probability distribution on $[m]\times[m]$, and $\tau$ is a tracial state on $\mathcal{A}(B)$, then there is a tracial state $\widetilde{\tau}$ on $\mathcal{A}(B^{\prime})$ such that $\tau=\widetilde{\tau}\circ\alpha$, $\operatorname{def}(\widetilde{\tau};\mu_{\pi})=\operatorname{def}(\tau;\mu_{\pi})$, and $\widetilde{\tau}(\sigma_{i_{1}}(x_{1})\cdots\sigma_{i_{k}}(x_{k}))=0$ for all sequences $x_{1},\ldots,x_{k}$ in $\Gamma$ and indices $i_{1},\ldots,i_{k}\in[m]$ such that $x_{j}\in U_{i_{j}}$ for all $1\leq j\leq k$. If $\tau$ is finite-dimensional (resp. Connes-embeddable), then $\widetilde{\tau}$ is also finite-dimensional (resp. Connes-embeddable).
  3. (c)
    For any $1\leq i\leq m$, the set $\{\prod_{x\in S}x:S\subseteq U_{i},|S|<n/2\}$ of monomials in $U_{i}$ of degree less than $n/2$ is linearly independent in $\mathcal{A}(U_{i},E_{i})$.
@@
Define $f_{i}:\mathbb{Z}_{2}^{U_{i}}\to\mathbb{Z}_{2}^{V_{i}}$ for each $i\in[m]$ by $f_{i}(\phi)(x)=\phi(x(1))\cdots\phi(x(n))$ for $\phi\in\mathbb{Z}_{2}^{U_{i}}$ and $x\in V_{i}$. By definition, $\phi\in E_{i}$ if and only if $f_{i}(\phi)\in C_{i}$, so $f_{i}(E_{i})=C_{i}$. If $f_{i}(\phi)(x)\neq f_{j}(\psi)(x)$ for some $\phi\in\mathbb{Z}_{2}^{U_{i}}$, $\psi\in\mathbb{Z}_{2}^{U_{j}}$, and $x\in V_{i}\cap V_{j}$, then we must have $\phi(x(i))\neq\psi(x(i))$ for some $i$. Since
$\sigma_{i}(x(1)\cdots x(n))=\sum_{\phi\in\mathbb{Z}_{2}^{U_{i}}}f_{i}(\phi)(x)\Phi_{U_{i},\phi}$
for all $x\in V_{i}$, $i\in[m]$, the functions $f_{i}$ correspond to a classical homomorphism $\alpha:\mathcal{A}(B)\to\mathcal{A}(B^{\prime})$ with $\alpha(\sigma_{i}(x))=\sigma_{i}(x(1)\cdots x(n))$ for all $i\in[m]$ and $x\in V_{i}$. This proves part (a).
Conversely, given $y\in\mathbb{Z}_{2}^{X\times[n-1]}$ and $\phi\in\mathbb{Z}_{2}^{V_{i}}$, define $\phi_{y}\in\mathbb{Z}_{2}^{U_{i}}$ by $\phi_{y}(x(1))=\phi(x)y(x,1)$, $\phi_{y}(x(j))=y(x,j-1)y(x,j)$ for $12\leq j\leq n-$ 1, and $\phi_{y}(x(n))=y(x,n-1)$. Since $f_{i}(\phi_{y})=\phi$, the function $\phi\mapsto\phi_{y}$ sends $C_{i}$ to $E_{i}$. Also if $\phi\in\mathbb{Z}_{2}^{V_{i}}$ and $\psi\in\mathbb{Z}_{2}^{V_{j}}$, then $\phi_{y}|_{U_{i}\cap U_{j}}\neq\psi_{y}|_{U_{i}\cap U_{j}}$ if and only if $\phi|_{V_{i}\cap V_{j}}\neq\psi|_{V_{i}\cap V_{j}}$, so the functions $\phi\mapsto\phi_{y}$ determine a classical homomorphism $\beta_{y}:\mathcal{A}(B^{\prime})\to\mathcal{A}(B)$ with $\beta_{y}(\sigma_{i}(x(1)))=\sigma_{i}(x)y(x,1)$, $\beta_{y}(\sigma_{i}(x(j)))=y(x,j-1)y(x,j)$ for $12\leq j\leq n-$ 1, and $\beta_{y}(\sigma_{i}(x(n)))=y(x,n-1)$ for all $i\in[m]$ and $x\in V_{i}$.
Given a tracial state $\tau$ on $\mathcal{A}(B)$, define a tracial state $\widetilde{\tau}$ on $\mathcal{A}(B^{\prime})$ by $\widetilde{\tau}=2^{-|X|(n-1)}\sum_{y}\tau\circ\beta_{y}$, where the sum is over all $y\in\mathbb{Z}_{2}^{X\times[n-1]}$. Notice that if $\tau$ is finite-dimensional (resp. Connes-embeddable), then $\widetilde{\tau}$ is also finite-dimensional (resp. Connes-embeddable). Since $\beta_{y}\circ\alpha$ is the identity on $\mathcal{A}(B)$, $\widetilde{\tau}\circ\alpha=\tau$. Since $\beta_{y}$ and $\alpha$ are $1$-homomorphisms,
@@
Finally, if $x_{1},\ldots,x_{k}$ is a sequence in $Z$, and $i_{1},\ldots,i_{k}$ is a sequence in $[m]$ such that $x_{j}\in U_{i_{j}}$, then there is an element $a\in\mathcal{A}(B)$ and set $S\subseteq X\times[n-1]$ such that
$\beta_{y}(\sigma_{i_{1}}(x_{1})\cdots\sigma_{i_{k}}(x_{k}))=m_{y}\tau(a)$
for all $y\in\mathbb{Z}_{2}^{X\times[n-1]}$, where $m_{y}:=\prod_{(x,j)\in S}y(x,j)$. If $x_{1},\ldots,x_{k}$ is in $\Gamma$, then $S$ is non-empty, and $\sum_{y}m_{y}=0$. Hence
$\widetilde{\tau}(\sigma_{i_{1}}(x_{1})\cdots\sigma_{i_{k}}(x_{k}))=2^{-|X|(n-1)}\sum_{y}m_{y}\tau(a)=0.$
This proves part (b).
For part (c), pick a tracial state $\tau$ on the finite-dimensional $C^{*}$-algebra $\mathcal{A}(V_{i},C_{i})$ (since $C_{i}$ is non-empty, this algebra is non-trivial). As in the proof of part (b), we can define a tracial state $\widetilde{\tau}=2^{|X|(n-1)}\sum_{y}\tau\circ\beta_{y}$ on $\mathcal{A}(U_{i},E_{i})$ with the property that $\widetilde{\tau}(x_{1}\cdots x_{k})=0$ if $11\leq k\leq n-$ 1 and $x_{1},\ldots,x_{k}\in U_{i}$ are distinct. If $S,T\subseteq U_{i}$, then
$\prod_{x\in S}x\cdot\prod_{x\in T}x=\prod_{x\in S\Delta T}x,$
where $S\Delta T:=(S\cup T)\setminus(S\cap T)$. If $|S|,|T|<n/2$, then $|S\Delta T|<n$, and $S\Delta T=\emptyset$ if and only if $S=T$. Hence by part (b),
$\widetilde{\tau}(\prod_{x\in S}x\cdot\prod_{x\in T}x)=\begin{cases}1&S=T\\ 0&S\neq T\end{cases}.$
It follows that the monomials $\{\prod_{x\in S}x:S\subseteq U_{i},|S|<n/2\}$ are linearly independent. ∎
@@
$Y=X\sqcup\{T_{i}(p,q),r_{i}(j,k):(i,p,q,j,k)\in{[m]\times[4]\times[d_{i}]\times[3]\times[d_{i}-1]}\}$
@@
Compared to [DFK+92], we’ve added the trivial constraints (4), as well as an extra row of the tableau. As mentioned above, the product in the constraints on the permutation-valued variables in parts (1)-(4) of the definition is the group product in $S_{5}$. The constraints in part (1) involve both original variables $x_{q}$ and permutation-valued variables $T_{i}(1,q)$, and say that the value of $T_{i}(1,q)$ is either $\pi^{(iq)}_{1}$ or $\pi^{(iq)}_{-1}$ depending on the value of $x_{q}$. In part (4), $x$ and $y$ can be either an original or a permutation-valued variable. If one of them is a permutation-valued variable, then all the corresponding boolean variables encoding the permutation-valued variable are included in the constraint (so the constraint on $x$ and $y$ may involve up to $14$ boolean variables). Since the constraints in part (4) are trivial, they do not contribute to $D_{i}$, but they are included in the list of clauses $(W_{ij},D_{ij})$ of the subdivided tableau. The point of the constraints in part (4) is that, with them, $\operatorname{Tab}_{sub}(B)$ is a subdivision of $\operatorname{Tab}(B)$. The extra row of the tableau is needed to compensate for the inclusion of these constraints in $\operatorname{Tab}_{sub}(B)$ (see 9.11). As in [DFK+92], the constraints $D_{i}$ encode the constraints $C_{i}$ as follows:
@@
Suppose $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, and let $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$. If $\psi\in D_{i}$, then $\psi|_{V_{i}}\in C_{i}$. Conversely, if $r\in S_{5}^{R_{i}}$, where $R_{i}=\{r_{i}(j,k):(j,k)\in[3]\times[d_{i}]\}$ is the set of randomizers in $W_{i}$, and $\phi\in C_{i}$, then there is a unique element $\phi_{r}\in D_{i}$ such that $\phi_{r}|_{V_{i}}=\phi$ and $\phi_{r}|_{R_{i}}=r$.
In this lemma, the statement that $\phi_{r}|_{R_{i}}=r$ means that for every randomizer $r_{i}(j,k)\in R_{i}$, the restriction of $\phi$ to the boolean variables corresponding to $r_{i}(j,k)$ is the encoding of the permutation $r(r_{i}(j,k))$.
Proof.
If $\psi\in D_{i}$, then by constraint (2), $\prod_{q}T_{i}(p+1,q)=\prod_{q}T_{i}(p,q)$. Since $\prod_{q}T_{i}(4,q)=\sigma_{i}$ by constraint (3), $\prod_{q}\pi_{x_{q}}^{(iq)}=\sigma_{i}$. Since the permutation branching program $P_{i}$ recognizes $C_{i}$, we conclude that $\psi|_{V_{i}}\in C_{i}$.
Conversely, given an assignment $r\in S_{5}^{R_{i}}$ to the variables $R_{i}$ and $\phi\in C_{i}$, we can set $T_{i}(1,q)=\pi_{\phi(x_{q})}^{(iq)}$ and $T_{i}(p+1,q)=r_{i}(p,q-1)^{-1}T_{i}(p,q)r_{i}(p,q)$ to get an assignment where $\prod_{q}T_{i}(4,q)=\sigma_{i}$. ∎
Although the permutation-valued variables in $\operatorname{Tab}(B)$ are shorthand for boolean variables, it is helpful to be able to work with the permutation-valued variables directly in $\mathcal{A}(\operatorname{Tab}(B))$. Suppose for a moment that $x_{1},\ldots,x_{7}$ are variables in a set $V$, and $C$ is a constraint on $V$ which includes the requirement that $x_{1},\ldots,x_{7}$ encode a permutation-valued variable $x$. Let $S=\{x_{1},\ldots,x_{7}\}$. If $\phi\in\mathbb{Z}_{2}^{S}$, then $\Phi_{S,\phi}=0$ in $\mathcal{A}(V,C)$ unless $\phi$ is the binary representation of an index $1200\leq j<$ 120, in which case we also write $\Phi_{S,\phi}$ as $\Phi_{S,j}$. Hence the subalgebra of $\mathcal{A}(V,C)$ is generated by the single unitary $\sum_{j=0}^{119}e^{2\pi ij/120}\Phi_{S,j}$, which we denote by the same symbol as the permutation-valued variable $x$. In particular, if $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ and $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$ as in Definition 9.6, then we can refer to $T_{i}(p,q)$ and $r_{i}(j,k)$ as unitary elements of $\mathcal{A}(W_{i},D_{i})$ of order $120$, and they generate the same subalgebra as the boolean variables encoding them. Since these variables do not occur in any other context $W_{j}$ for $j\neq i$, we also use $T_{i}(p,q)$ and $r_{i}(j,k)$ to refer to $\sigma_{i}(T_{i}(p,q))$ and $\sigma_{i}(r_{i}(j,k))$ in $\mathcal{A}(\operatorname{Tab}(B))$. We use the same convention for $\mathcal{A}(W_{i\ell},D_{i\ell})$, although since the variables $T_{i}(p,q)$ and $r_{i}(j,k)$ occur in more than one constraint of $\operatorname{Tab}_{sub}(B)$, we are stuck with the notation $\sigma_{i\ell}(T_{i}(p,q))$ and $\sigma_{i\ell}(r_{i}(j,k))$ when refering to these variables in $\mathcal{A}(\operatorname{Tab}_{sub}(B))$. With these conventions, we can state the following noncommutative version of Lemma 9.7.
@@
Suppose that $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, and let $\operatorname{Tab}(B)=(Y,\{(W_{i},D_{i})\}_{i=1}^{m})$. Let $R_{i}=\{r_{i}(j,k):(j,k)\in[3]\times[d_{i}-1]\}$ be the set of randomizers in $W_{i}$, and let $R=\bigcup_{i}R_{i}$.
  1. (a)
    The natural map
    $\mathcal{A}(V_{i},C_{i})\otimes\mathbb{C}\mathbb{Z}_{120}^{R_{i}}\to\mathcal{A}(W_{i},D_{i}):x_{i}\mapsto x_{i},r_{i}(j,k)\mapsto r_{i}(j,k)$
    is an isomorphism. In particular, $\mathcal{A}(W_{i},D_{i})$ is generated as an algebra by $V_{i}\cup R_{i}$, and $\mathcal{A}(\operatorname{Tab}(B))$ is generated by $\bigcup_{i}\{\sigma_{i}(x):x\in V_{i}\}\cup R$.
  2. (b)
    The natural inclusion $\alpha:\mathcal{A}(B)\to\mathcal{A}(\operatorname{Tab}(B))$ defined by $\alpha(\sigma_{i}(x))=\sigma_{i}(x)$ for $i\in[m]$ and $x\in V_{i}$ is a classical homomorphism.
  3. (c)
    If $r\in S_{5}^{R}$, then there is a classical homomorphism $\beta_{r}:\mathcal{A}(\operatorname{Tab}(B))\to\mathcal{A}(B)$ such that for all $i\in[m]$, if $x\in V_{i}$ then $\beta_{r}(\sigma_{i}(x))=\sigma_{i}(x)$, and if $x\in R_{i}$ then $\beta_{r}(x)=e^{2\pi ij/120}$ where $r(x)=\gamma_{j}$ in the enumeration of $S_{5}$ fixed above.
  4. (d)
    Let $\mathcal{M}$ be the set of monomials in $\mathcal{A}(B)$ of the form $u\sigma_{i}(z)^{a}v$, where $z\in\mathcal{R}_{i}$ for some $i\in[m]$, $1201\leq a<$ 120, and $u$ and $v$ are monomials in $\{\sigma_{j}(x):j\in[m],x\in V_{j}\cup R_{j}\}$ which do not contain $z$. If $\pi$ is a probability distribution on $[m]\times[m]$, and $\tau$ is a tracial state on $\mathcal{A}(B)$, then there is a tracial state $\widetilde{\tau}$ on $\mathcal{A}(\operatorname{Tab}(B))$ such that $\tau=\widetilde{\tau}\circ\alpha$, where $\alpha$ is the classical homomorphism from part (b), $\operatorname{def}(\widetilde{\tau};\mu_{\pi})=\operatorname{def}(\tau;\mu_{\pi})$, and $\widetilde{\tau}(y)=0$ for all $y\in\mathcal{M}$. Furthermore, if $\tau$ is finite-dimensional (resp. Connes-embeddable), then $\widetilde{\tau}$ is also finite-dimensional (resp. Connes-embeddable).
Proof.
For part (a), the algebra $\mathbb{C}\mathbb{Z}_{120}^{R_{i}}$ has a basis consisting of the joint spectral projections
$\Phi_{R_{i},r}=\prod_{x\in R_{i}}\lambda(r(x))^{-1}\prod_{\gamma_{k}\neq r(x)}(x-e^{2\pi ik/120}),\quad r\in\mathbb{Z}_{120}^{R_{i}},$
where $\lambda(\gamma_{j})=\prod_{k\neq j}(e^{2\pi ij/120}-e^{2\pi ik/120})$. Hence $\mathcal{A}(V_{i},C_{i})\otimes\mathbb{C}\mathbb{Z}_{120}^{R_{i}}$ has a basis consisting of the elements $\Phi_{V_{i},\phi}\otimes\Phi_{R_{i},r}$ for $\phi\in C_{i}$ and $r\in\mathbb{Z}_{120}^{R_{i}}$. Using the enumeration of $S_{5}$ fixed earlier, we can interpret $\mathbb{Z}_{120}^{R_{i}}$ as the set $S_{5}^{R_{i}}$ of permutation-valued assignments to $R_{i}$. The natural homomorphism $\mathcal{A}(V_{i},C_{i})\otimes\mathbb{C}\mathbb{Z}_{120}^{R_{i}}\to\mathcal{A}(W_{i},D_{i})$ sends $\Phi_{V_{i},\phi}\otimes\Phi_{R_{i},r}$ to $\sum_{\psi}\Phi_{W_{i},\psi}$, where the sum is across all $\psi\in D_{i}$ such that $\psi|_{V_{i}}=\phi$ and $\psi|_{R_{i}}=r$. By Lemma 9.7, the restriction map $\phi\mapsto\phi|_{V_{i}\cup R_{i}}$ is a bijection between $D_{i}$ and $C_{i}\times S_{5}^{R_{i}}$, so this homomorphism is an isomorphism.
Parts (b) and part (c) follow immediately from Lemma 9.7 and the definition of a classical homomorphism. Alternatively, part (b) also follows from Corollary 6.7.
The proof of part (d) is similar to the proof of Lemma 9.4, part (b). Given a tracial state $\tau$ on $\mathcal{A}(B)$, let $\widetilde{\tau}$ be the tracial state on $\mathcal{A}(\operatorname{Tab}(B))$ defined by $\widetilde{\tau}=\tfrac{1}{120^{|R|}}\sum_{r}\tau\circ\beta_{r}$, where the sum is over $r\in S_{5}^{R}$. If $\tau$ is finite-dimensional (resp. Connes-embeddable), then $\widetilde{\tau}$ is finite-dimensional (resp. Connes-embeddable). Since $\beta_{r}(\sigma_{i}(x))=\sigma_{i}(x)$ for all $i\in[m]$ and $x\in V_{i}$, $\beta_{r}\circ\alpha$ is the identity on $\mathcal{A}(B)$, and $\widetilde{\tau}\circ\alpha=\tau$. By parts (b) and (c), $\operatorname{def}(\tau\circ\beta_{r})\leq\operatorname{def}(\tau)=\operatorname{def}(\tau\circ\beta_{r}\circ\alpha)\leq\operatorname{def}(\tau\circ\beta_{r})$. This means that $\operatorname{def}(\tau\circ\beta_{r})=\operatorname{def}(\tau)$, so $\operatorname{def}(\tau)=\operatorname{def}(\widetilde{\tau})$. Finally, suppose $y\in\mathcal{M}$, so $y=u\sigma_{i}(z)^{a}v$ for some $z\in R_{i}$, $1201\leq a<$ 120, and monomials $u$, $v$ which do not contain $z$. By part (c), there is some monomial $y^{\prime}$ in $\{\sigma_{j}(x):j\in[m],x\in V_{i}\}$ such that for all $r\in S_{5}^{R}$, we have $\beta_{r}(y)=e^{2\pi aij/120}c_{r^{\prime}}y^{\prime}$, where $r(z)=\gamma_{j}$, and $c_{r^{\prime}}\in\mathbb{C}$ depends only on $r^{\prime}=r|_{R\setminus\{z\}}$. Hence
$\widetilde{\tau}(y)=\tfrac{1}{120^{|R|}}\sum_{r\in S_{5}}\tau(\beta_{r}(y))=\tfrac{1}{120^{|R|}}\sum_{j=0}^{120}e^{2\pi aij/120}\sum_{r^{\prime}\in S_{5}^{R\setminus\{z\}}}c_{r^{\prime}}\tau(y^{\prime})=0,$
finishing the proof of part (d). ∎
We need one more general fact about permutation-valued variables.